# Conformal Prediction Adaptive to Unknown Subpopulation Shifts

### v1

@@ 3.2 Conformal Prediction with multicalibrated domain classifier
In most cases, training a perfect classifier is impossible. Therefore, since $c$ can only provide the estimated probability distribution, how well calibrated $c$ is matters a lot to the coverage provided by Algorithm 1, especially in cases where the in-domain distributions differ a lot. Therefore, it’s more feasible to train a domain classifier that makes mistakes within a limited range. We will use the notion of multicalibration, which is used to measure fairness of a predictor DBLP:journals/corr/abs-1711-08513 .
@@
Denote $\mathcal{G}$ as a family of subsets of $\mathcal{X}$, $\tilde{p}:\mathcal{X}\rightarrow\Delta^{K}$ as some predictor, and $p^{*}:\mathcal{X}\rightarrow\Delta^{K}$ as the true probability distribution. $\tilde{p}$ is multicalibrated with respect to $\mathcal{G}$ if for all $v\in\tilde{p}(\mathcal{X})$ and $G\in\mathcal{G}$,
$\mathbb{E}(p^{*}(x)|x\in G,\tilde{p}(x)=v)=v.$
By defining the family of subsets, $\mathcal{G}$, as the set of all possible test environments and assuming that the domain classifier, $c$, from Algorithm 1 is multicalibrated with respect to $\mathcal{G}$, we can ensure coverage conditioned on each test environment as shown in Theorem 3.3
@@
Suppose $\mathcal{G}$ is the set of all test environments and $c$ is multicalibrated with respect to $\mathcal{G}$. Then the output of Algorithm 1, $\mathcal{C}_{\alpha}$, satisfies
$Pr(Y_{\text{test}}\in C(X_{\text{test}})|(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{\text{test}})\geq 1-\alpha$
@@ 3.3 Conformal Prediction with multiaccurate domain classifier
While learning multicalibrated predictors is easier than learning the Bayes-optimal classifier, they are still shown to have high computational and sample complexity which makes it difficult to train gopalan2022lowdegreemulticalibration . Therefore, an even more relaxed assumption is necessary in most cases, which motivates us to use the notion of multiaccuracy DBLP:journals/corr/abs-1805-12317 .
@@
Denote $\mathcal{G}$ as a family of subsets of $\mathcal{X}$, $\tilde{p}:\mathcal{X}\rightarrow\Delta^{K}$ as some predictor, and $p^{*}:\mathcal{X}\rightarrow\Delta^{K}$ as the true probability distribution. $\tilde{p}$ is multiaccurate with respect to $\mathcal{G}$ if for all $G\in\mathcal{G}$,
$\mathbb{E}(p^{*}(x)|x\in G)=\mathbb{E}(\tilde{p}(x)|x\in G).$
Under Definition 3.4, multiaccuracy relaxes the definition of multicalibration and only requires a predictor to be calibrated within a subset of $\mathcal{X}$. Therefore, we propose Algorithm 2, where $\hat{\lambda}$ is the mean of $c(X^{i}_{\text{test}})$ over the test data set instead. By defining the family of subsets, $\mathcal{G}$, as the set of all possible test environments and assuming that $c$ from Algorithm 2 is multiaccurate, we can ensure coverage conditioned on each test environment as shown in Theorem 3.5.
@@
Suppose $\mathcal{G}$ is the set of all test environments and $c$ is multiaccurate with respect to $\mathcal{G}$. Then the output of Algorithm 2, $\mathcal{C}_{\alpha}$, satisfies
$Pr(Y_{\text{test}}\in C(X_{\text{test}})|(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{test})\geq 1-\alpha$
@@
As mentioned previously, multicalibration is a property that is difficult to formally prove. However, hansen2024multicalibrationpostprocessingnecessary conducted a comprehensive study evaluating the effectiveness of multicalibration post-processing across diverse datasets (including tabular, image, and language) and models ranging from simple decision trees to 90 million parameter fine-tuned LLMs. The study found that models which are well trained tend to be relatively multicalibrated without requiring additional post-processing. Thus, we believe that assuming access to a multi-accuracy classifier $c$ is an easy to satisfy assumption.
4 Adapting to subpopulation shifts without domain information
@@ 4.1 Conformal Prediction weighted by similarity measures
In many real word tasks, similarity measures in the representation space often capture the semantic similarity between images or languages. Therefore, we propose Algorithm 3 which assumes that data with higher similarities in the embedding space have higher probability to be from the same domain. Algorithm 3 is exactly the weighted conformal prediction method proposed by tibshirani2020conformalpredictioncovariateshift where instead of weighing the calibration data by the likelihood ratio, we propose weighing the calibration data by similarity between the embedding of each calibration data and the test data. Weighting by similarity measures assumes that data with high similarity measures are semantically similar, i.e., from the same or similar domains. However, empirical results show that such assumption is not true across all domains, therefore, we propose keeping only a fraction of the data with the highest similarity measures to the test data. The percentage of data to include is defined as $\beta$ in Algorithm 3.
@@ 4.2 Conformal risk control for LLM hallucination detection
where $Y^{*}_{\text{test}}$ is the greedy output to the query $X_{\text{test}}$, $Y_{\text{test}}$ is the ground truth and $A(X_{\text{test}},Y_{\text{test}},Y^{*}_{\text{test}})=1$ if $Y^{*}_{\text{test}}$ is a hallucinated response to query $X_{\text{test}}$ and 00 otherwise. We will follow the steps from Algorithm 3 and make necessary adjustments. Specifically, first, since we wish to bound the recall error, all calibration data are hallucinated generations. Second, we compute the score using score function $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, which uses a generative model $\hat{f}$. We note that this scoring function is different from the score functions from the vision tasks, as the score does not take the ground truth into account. We then follow the same steps in Algorithm 3 to find the threshold $\hat{q}_{\alpha}$ where we let $\alpha$ to be $1-r_{\text{test}}$. Lastly, we label the test data “hallucination” if the score is above $\hat{q}_{\alpha}$ and “not hallucination” otherwise.
@@
$\hat{q}_{\alpha}\leftarrow\text{Quantile}\left(1-\alpha,\sum^{n^{\prime}+1}_{i=1}m_{i}\delta_{s_{i}}\right)$triangleright $\delta_{s_{i}}$ denotes a point mass at $s_{i}$
@@
5 Experiments
Figure 2: Coverage with 26 domains and 3 classes per domain. Vision transformer is calibrated with LAC score function for various algorithms. (Left) Coverage across 100 test environments at $\alpha=0.05$. Each coverage data is the average of 15 calibration/test splits. Mean and standard deviations are shown in the legend. (Right) Mean and standard deviation of coverage across 100 test environments. Our algorithms (A1, A2, and oracle) demonstrate the desired coverage of 0.95 across test environments (unlike unweighted) and with minimal over-coverage (unlike max). Further, the practical algorithms A1 and A2 quite closely match the ideal oracle coverage.
@@ 5.1 Experimental Setup for vision tasks
Dataset.
For the vision tasks, we use the ImageNet Large Scale Visual Recognition Challenge dataset ILSVRC15 , which contains 1000 classes. We split the validation data in two, half as the calibration set and the other half as the test set. The split is done multiple times as the coverage guarantee of conformal prediction is over the randomness of the calibration set. To simulate subpopulation shifts, we adopt the BREEDS methodology santurkar2020breedsbenchmarkssubpopulationshift . The method creates a tree structure where the leaf nodes are the 1000 classes and the internal nodes are superclasses. We picked the nodes at level 3 as our domains and the descendents of each node are the classes in each domain. To create a balanced train environment, we keep the number of classes in each domain the same by removing domains with non-sufficient number of classes and removing some classes from domains with too many classes. We test on two different number of classes, one with 26 domains with 3 classes each and the other with 15 domains and 17 classes each. To simulate the different test environments, we follow the sampling strategy from hsu2019measuringeffectsnonidenticaldata to draw $\lambda$ from a Dirichlet distribution with parameter $\alpha^{\prime}$. The parameter $\alpha^{\prime}$ controls the heterogeneity, i.e., as $\alpha^{\prime}\rightarrow 0$, $\lambda$ is 1 for one domain and 0 for all others. As $\alpha^{\prime}\rightarrow\infty$, $\lambda$ becomes uniform which reduces the problem to the no subpopulation shift case. Code is available at https://github.com/mrreganwang/Conformal_Subpop_Shifts
Models.
We test on three different pretrained models: resnet50 pretrained on ImageNet he2015deepresiduallearningimage , vision transformer pretrained on ImageNet21k and finetuned on ImageNet 2021 steiner2021augreg ; dosovitskiy2020vit ; rw2019timm , and vision transformer pretrained on WIT-400M image-text pairs by OpenAI using CLIP embedding and finetuned on ImageNet-1k Radford2021LearningTV ; cherti2022reproducible ; dosovitskiy2020vit ; rw2019timm . For the domain classifiers, we modified the fully-connected layers of the three pre-trained models. The modified fully-connected layers now includes three dense layers with sizes 2048, 1024, and 512. The output layer is a softmax layer with output size of either 26 or 15.
Domain Classifier Training.
For training, only the last 3 fully connected layers are updated. The training uses Adam kingma2017adammethodstochasticoptimization with cross entropy loss. After training, the domain classifiers are then calibrated using Multi-domain temperature scaling introduced in yu2022robustcalibrationmultidomaintemperature to reduce calibration error.
@@ 5.2 Main Results
Coverage with varying test environments.
We calibrated a pre-trained vision transformer with LAC score function and tested it on test set sampled from 100 different test environments. The test environment consists of 26 domains, with 3 classes in each domain while the $\lambda$ was sampled from a Dirichlet distribution with parameter 0.1. Each coverage datapoint is averaged across 15 random calibration/test split. The results are plotted in Figure  2. From Figure  2 (Left) we observe that all three proposed algorithms were able to provide coverage for all test environments while standard conformal prediction could not for some test environments. For the max method, which conformalize the model using the worst case method mentioned in section 2.2, we see that marginal coverage is satisfied for all test environments, however, they are severely over-covered. We also observe that when compared to the standard conformal prediction, the standard deviations for the proposed algorithms are much smaller. This shows the adaptiveness of the proposed algorithms to maintain the desired coverage across test environments. From Figure 2 (Right), we see that the proposed algorithms are able to maintain coverage, while ensuring low standard deviations across different $1-\alpha$.
Coverage under different settings
We obtained the coverage results with varying score functions, model architectures, and degree of subpopulation shifts which we present in Appendix C. The coverage results are consistent across different settings.
@@ 6.1 Experimental Setup for language tasks
Datasets
To simulate different domains in generative language tasks, we use two distinct datasets: TriviaQA joshi2017triviaqa , a closed-book question answering dataset, and GSM8K cobbe2021gsm8k , a mathematical reasoning benchmark. Specifically, we use 2,500 samples from the test split of TriviaQA and the full GSM8K test set, which contains 1,319 questions. To create the calibration and test data, we first randomly select 500 TriviaQA samples and 500 GSM8K samples to create the test set. The rest of the samples are used as the calibration set. To keep the calibration set balanced, we randomly removed 1181 TriviaQA samples, resulting in a calibration set with 1638 samples. We repeat this process 10 times. To simulate each test environment, we again draw $\lambda$ from a dirichlet distribution with parameter 0.1 and remove test data from each of the two domains to match the $\lambda$.
Models
We use LLaMA-3-8B llama3modelcard as the generative model and obtain responses via greedy decoding. Following prior work lin2023generating ; bakman2024mars , we employ GPT-4o openai2023gpt4 as the correctness evaluator, using the query, generated response, and ground truth answer(s) as input. To assess the similarity between test samples and calibration data points, we use the all-mpnet-base-v2 model from SentenceTransformers reimers-2019-sentence-bert .
Figure 3: Coverage with 26 domains and 3 classes per domain. Vision transformer is calibrated with LAC score function for various algorithms. For the results of Algorithm 3, the parameters $\sigma$ and $\beta$ are 0.7 and $0.1$ respectively. (Left) Coverage across 100 test environments at $\alpha=0.05$. Each coverage data is the average of 15 calibration/test splits. Mean and standard deviations are shown in the legend. (Right) Mean and standard deviation of coverage across 100 test environments. Our algorithm (A3 in pink) demonstrates the desired coverage of 0.95 across test environments with minimal over-coverage. Further, even without using any distributional or domain information, it matches the ideal coverage of the oracle (in green) which knows the test distribution exactly.

### v2

@@ 3.2 Conformal Prediction with multicalibrated domain classifier
In most cases, training a perfect classifier is impossible. Therefore, since $c$ can only provide the estimated probability distribution, how well calibrated $c$ is matters to the coverage provided by Algorithm 1, especially in cases where the in-domain distributions differ a lot. Therefore, it’s more feasible to train a domain classifier that makes mistakes within a limited range. We will use the notion of multicalibration, which is used to measure fairness of a predictor (Hébert-Johnson et al., 2017).
@@
[Multicalibrated domain classifier] Denote $\mathcal{D}$ as a family of distributions on $\mathcal{X}$, $c:\mathcal{X}\rightarrow\Delta^{K}$ as the trained domain classifier, and $c^{*}:\mathcal{X}\rightarrow\Delta^{K}$ as the perfect domain classifier. $c$ is multicalibrated with respect to $\mathcal{D}$ if for all $v\in c(\mathcal{X})$ and $D\in\mathcal{D}$,
$\mathbb{E}(c^{*}(x)|x\sim D,c(x)=v)=v.$
By defining $\mathcal{D}$, as the set of all possible test environments and assuming that the domain classifier, $c$, from Algorithm 1 is multicalibrated with respect to $\mathcal{D}$, we can ensure coverage conditioned on each test environment as shown in Theorem 3.3
@@
[Multicalibrated domain classifier implies coverage under subpopulation shifts] Suppose $c:\mathcal{X}\rightarrow\Delta^{K}$ is a domain classifier that maps the input to a probability distribution over the $K$ domains and $(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{\text{test}}$, as defined in Section 2.1. Furthermore, suppose $\mathcal{D}$ is the set of all possible $\mathbb{P}_{\text{test}}$ and $c$ is multicalibrated with respect to $\mathcal{D}$, as defined in Definition 3.2. Then the output of Algorithm 1, $\mathcal{C}_{\alpha}$, satisfies
$Pr(Y_{\text{test}}\in C(X_{\text{test}}))\geq 1-\alpha.$
@@ 3.3 Conformal Prediction with multiaccurate domain classifier
While learning multicalibrated predictors is easier than learning the Bayes-optimal classifier, they are still shown to have high computational and sample complexity which makes it difficult to train (Gopalan et al., 2022). Therefore, an even more relaxed assumption is necessary in most cases, which motivates us to use the notion of multiaccuracy (Kim et al., 2018).
@@
[Multiaccurate domain classifier] Denote $\mathcal{D}$ as a family of distributions on $\mathcal{X}$, $c:\mathcal{X}\rightarrow\Delta^{K}$ as the trained domain classifier, and $c^{*}:\mathcal{X}\rightarrow\Delta^{K}$ as the perfect domain classifier. $c$ is multiaccurate with respect to $\mathcal{D}$ if for all $D\in\mathcal{D}$,
$\mathbb{E}(c^{*}(x)|x\sim D)=\mathbb{E}(c(x)|x\sim D).$
Under Definition 3.4, multiaccuracy relaxes the definition of multicalibration and only requires a predictor to be calibrated within a subset of $\mathcal{X}$. Therefore, we propose Algorithm 2, where $\hat{\lambda}$ is the mean of $c(X^{i}_{\text{test}})$ over the test data set instead. By defining the family of subsets, $\mathcal{D}$, as the set of all possible test environments and assuming that $c$ from Algorithm 2 is multiaccurate, we can ensure coverage conditioned on each test environment as shown in Theorem 3.5.
@@
[Multiaccurate domain classifier implies coverage under subpopulation shifts] Suppose $c:\mathcal{X}\rightarrow\Delta^{K}$ is a domain classifier that maps the input to a probability distribution over the $K$ domains and $(X_{\text{test}},Y_{\text{test}})\sim\mathbb{P}_{\text{test}}$, as defined in section 2.1. Furthermore, suppose $\mathcal{D}$ is the set of all possible $\mathbb{P}_{\text{test}}$ and $c$ is multiaccurate with respect to $\mathcal{D}$, as defined in Definition 3.4. Then the output of Algorithm 2, $\mathcal{C}_{\alpha}$, satisfies
$Pr(Y_{\text{test}}\in C(X_{\text{test}}))\geq 1-\alpha.$
@@
Multicalibration is difficult to prove formally. However, Hansen et al. (2024) conducted a comprehensive study and show that well trained models tend to be relatively multicalibrated. Thus, we believe that assuming access to a multi-accuracy classifier $c$ (significantly easier to satisfy than multi-calibration) is an easy to satisfy assumption.
4 Subpopulation shifts without any domain classifier
@@ 4.1 Conformal Prediction weighted by similarity measures
In many real word tasks, similarity measures in the representation space often capture the semantic similarity between images or languages. Therefore, we propose Algorithm 3 which assumes that data with higher similarities in the embedding space have higher probability to be from the same domain. Algorithm 3 is exactly the weighted conformal prediction method proposed by Tibshirani et al. (2020) where instead of weighing the calibration data by the likelihood ratio, we propose weighing the calibration data by similarity between the embedding of each calibration data and the test data. Weighting by similarity measures assumes that data with high similarity measures are semantically similar, i.e., from the same or similar domains. However, empirical results show that such assumption is not true across all domains, therefore, we propose keeping only a fraction of the data with the highest similarity measures to the test data. The percentage of data to include is defined as $\beta$ in Algorithm 3.
@@ 4.2 Conformal risk control for LLM hallucination detection
where $Y^{*}_{\text{test}}$ is the greedy output to the query $X_{\text{test}}$, $Y_{\text{test}}$ is the ground truth and $A(X_{\text{test}},Y_{\text{test}},Y^{*}_{\text{test}})=1$ if $Y^{*}_{\text{test}}$ is a hallucinated response to query $X_{\text{test}}$ and 0 otherwise. We will follow the steps from Algorithm 3 and make necessary adjustments. Specifically, first, since we wish to bound the recall error, all calibration data are hallucinated generations. Second, we compute the score using score function $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, which uses a generative model $\hat{f}$. We note that this scoring function is different from the score functions from the vision tasks, as the score does not take the ground truth into account. We then follow the same steps in Algorithm 3 to find the threshold $\hat{q}_{\alpha}$ where we let $\alpha$ to be $1-r_{\text{test}}$. Lastly, we label the test data “hallucination” if the score is above $\hat{q}_{\alpha}$ and “not hallucination” otherwise.
@@
$\hat{q}_{\alpha}\leftarrow\text{Quantile}\left(1-\alpha,\sum^{n^{\prime}+1}_{i=1}m_{i}\delta_{s_{i}}\right)$ $\triangleright$ $\delta_{s_{i}}$ denotes a point mass at $s_{i}$
@@
5 Experiments with knowledge of domains
Figure 2: Coverage distribution over 100 test environments with subpopulation shifts. (Left) Coverage across 100 test environments generated using Dirichlet sampling over 26 domains, and the averaged over 15 calibration/test splits. Mean and standard deviations are shown in the legend. (Right) Mean and standard deviation of coverage across 100 test environments. Note that max tends to substantially over cover compared to desired coverage of 0.95. Our algorithms (A1, A2, and oracle) demonstrate the desired coverage across test environments (unlike unweighted and Conditional Calibration that have significant under-coverage). They also have minimal over-coverage and tightly follow the target (unlike max which significantly over-covers). Further, the practical algorithms A1 and A2 quite closely match the ideal oracle coverage.
@@ 5.1 Experimental Setup for vision tasks
Dataset. For the vision tasks, we use the ImageNet Large Scale Visual Recognition Challenge dataset (Russakovsky et al., 2015), which contains 1000 classes. We split the validation data in two, half as the calibration set and the other half as the test set. The split is done multiple times as the coverage guarantee of conformal prediction is over the randomness of the calibration set. To simulate subpopulation shifts, we adopt the BREEDS methodology (Santurkar et al., 2020). The method creates a tree structure where the leaf nodes are the 1000 classes and the internal nodes are superclasses. We picked the nodes at level 3 as our domains and the descendents of each node are the classes in each domain. To create a balanced train environment, we keep the number of classes in each domain the same by removing domains with non-sufficient number of classes and removing some classes from domains with too many classes. We test on two different number of classes, one with 26 domains with 3 classes each and the other with 15 domains and 17 classes each. To simulate the different test environments, we follow the sampling strategy from Hsu et al. (2019) to draw $\lambda$ from a Dirichlet distribution with parameter $\alpha^{\prime}$. The parameter $\alpha^{\prime}$ controls the heterogeneity, i.e., as $\alpha^{\prime}\rightarrow 0$, $\lambda$ is 1 for one domain and 0 for all others. As $\alpha^{\prime}\rightarrow\infty$, $\lambda$ becomes uniform which reduces the problem to the no subpopulation shift case.
Models. We test on three different pretrained models: resnet50 pretrained on ImageNet (He et al., 2015), vision transformer pretrained on ImageNet21k and finetuned on ImageNet 2021 (Steiner et al., 2021; Dosovitskiy et al., 2021; Wightman, 2019), and vision transformer pretrained on WIT-400M image-text pairs by OpenAI using CLIP embedding and finetuned on ImageNet-1k (Radford et al., 2021; Cherti et al., 2022; Dosovitskiy et al., 2021; Wightman, 2019). For the domain classifiers, we modified the fully-connected layers of the three pre-trained models. The modified fully-connected layers now includes three dense layers with sizes 2048, 1024, and 512. The output layer is a softmax layer with output size of either 26 or 15.
Domain Classifier Training. For training, only the last 3 fully connected layers are updated. The training uses Adam (Kingma & Ba, 2017) with cross entropy loss. After training, the domain classifiers are then calibrated using Multi-domain temperature scaling introduced in Yu et al. (2022) to reduce calibration error.
@@ 5.2 Main Results
Coverage with varying test environments. We calibrated a pre-trained vision transformer with LAC score function and tested it on test set sampled from 100 different test environments. The test environment consists of 26 domains, with 3 classes in each domain while the $\lambda$ was sampled from a Dirichlet distribution with parameter 0.1. Each coverage datapoint is averaged across 15 random calibration/test split. The results are plotted in Figure  2. From Figure  2 (Left) we observe that all three proposed algorithms were able to provide coverage for all test environments while standard conformal prediction could not for some test environments. For the max method, which conformalize the model using the worst case method mentioned in section 2.1, we see that marginal coverage is satisfied for all test environments, however, they are severely over-covered. We also observe that when compared to the standard conformal prediction, the standard deviations for the proposed algorithms are much smaller. This shows the adaptiveness of the proposed algorithms to maintain the desired coverage across test environments. From Figure 2 (Right), we see that the proposed algorithms are able to maintain coverage, while ensuring low standard deviations across different $1-\alpha$.
Coverage under different settings. We obtained the coverage results with varying score functions, model architectures, and degree of subpopulation shifts which we present in Appendix D. The coverage results are consistent across different settings.
Figure 3: Adapting to subpopulation shifts without a domain classifier. Vision transformer is calibrated with LAC score function for various algorithms. For the results of Algorithm 3, the parameters $\sigma$ and $\beta$ are 0.7 and $0.1$ respectively. (Left) Coverage across 100 test environments at $\alpha=0.05$. Each coverage data is the average of 15 calibration/test splits. Mean and standard deviations are shown in the legend. (Right) Mean and standard deviation of coverage across 100 test environments. Our algorithm (A3 in pink) demonstrates the desired coverage of 0.95 across test environments with minimal over-coverage. Further, even without using any distributional or domain information, it matches the ideal coverage of the oracle (in green) which knows the test distribution exactly.
@@ 6.1 Experimental Setup for language tasks
Datasets. To simulate different domains in generative language tasks, we use two distinct datasets: TriviaQA (Joshi et al., 2017), a closed-book question answering dataset, and GSM8K (Cobbe et al., 2021), a mathematical reasoning benchmark. Specifically, we use 2,500 samples from the test split of TriviaQA and the full GSM8K test set, which contains 1,319 questions. To create the calibration and test data, we first randomly select 500 TriviaQA samples and 500 GSM8K samples to create the test set. The rest of the samples are used as the calibration set. To keep the calibration set balanced, we randomly removed 1181 TriviaQA samples, resulting in a calibration set with 1638 samples. We repeat this process 10 times. To simulate each test environment, we again draw $\lambda$ from a dirichlet distribution with parameter 0.1 and remove test data from each of the two domains to match the $\lambda$.
Models. We use LLaMA-3-8B (AI@Meta, 2024) as the generative model and obtain responses via greedy decoding. Following prior work (Lin et al., 2024; Bakman et al., 2024), we employ GPT-4o (OpenAI, 2023) as the correctness evaluator, using the query, generated response, and ground truth answer(s) as input. To assess the similarity between test samples and calibration data points, we use the all-mpnet-base-v2 model from SentenceTransformers (Reimers & Gurevych, 2019).